# Article Category

### v1

@@ Abstract
Deploying autonomous robots in safety-critical applications requires safety guarantees. Provably safe reinforcement learning is an active field of research which aims to provide such guarantees using safeguards. These safeguards should be integrated during training to prevent a large sim-to-real gap. While there are several approaches for safeguarding sampling-based reinforcement learning, analytic gradient-based reinforcement learning often achieves superior performance and sample efficiency. However, there is no safeguarding approach for this learning paradigm yet. Our work addresses this gap by developing the first effective safeguard for analytic gradient-based reinforcement learning. We analyse existing, differentiable safeguards, adapt them through modified mappings and gradient formulations, and integrate them with a state-of-the-art learning algorithm and a differentiable simulation. We evaluate how different safeguards affect policy optimisation using numerical experiments on two classical control tasks. The results demonstrate safeguarded training without compromising performance.
@@ 1 INTRODUCTION
The transfer of physical labour from humans and human-operated machines to robots is a long-standing goal of robotics research. While robots have been successfully deployed in controlled environments, such as factories, deploying strong robots in human proximity is still a challenge [StrongRobot, StrongRobot2]. One reason is a lack of safety guarantees to ensure that robots do not harm humans or themselves [SafetyRequired].
Classical control algorithms, such as PID or MPC controllers, can be designed for safety [SafePID, RobustMPC, MPC2]. While reinforcement learning controllers do not provide safety guarantees, they often achieve state-of-the-art performance [GreatRL, GreatRL2] due to their broad applicability, representation power, and fast inference. In recent years, provably safe reinforcement learning has emerged as a research field to provide safety guarantees [ProvablySafeSurvey, ProvablySafeRL].
Current safeguards are applied in conjunction with reinforcement learning algorithms that rely on the policy gradient theorem to estimate reward landscapes [ActionReplacement1, ActionProjection, RayMasking, BoundaryProjectionLoss, Gauge, BoundaryProjection]. With the advent of differentiable physics simulators [Mujoco, brax, ChainQueen, Thuerey, SAPO], analytically computing the gradient of the reward with respect to the actions has become possible. Reinforcement learning algorithms that exploit these gradients promise faster training and better performance [GradientRL1, GradientRL2, PODS, SHAC]. While these simulators require approximations to remain differentiable, maintaining simulation accuracy is possible. As a result, one can train the policy exclusively in simulation before deploying it to real systems and avoid costly real-world training [Real1, Real2].
@@
Applying safeguards during training is desirable for safety-critical systems to prevent an effectively enlarged sim-to-real gap [simtoreal1, simtoreal2]. Otherwise, the optimisation may converge to strategies that rely on unsafe states or actions. When deployment safeguards subsequently restrict these strategies, the policy can become suboptimal or even fail for non-convex objective landscapes, as it has not learned alternative, safe solutions [DuringTraining]. While crafting reward functions that reliably encode safety requirements could theoretically prevent performance degradation in deployment, this is notoriously difficult without introducing unintended incentives [RewardHacking1, RewardHacking2]. Moreover, safeguards can guide the exploration process to aid learning in challenging solution spaces [RayMasking, PostTraining]. However, there are currently no safeguarding attempts for analytic gradient-based reinforcement learning.
@@
Our work addresses this by combining state-of-the-art analytic gradient-based reinforcement learning algorithms, differentiable safeguards, and differentiable simulations [^1] Code available at https://github.com/TimWalter/SafeGBPO. To ensure safety, we incorporate a range of existing set-based safeguarding methods that map policy actions into a safe action set. We formalise desirable safeguarding properties in the context of differentiable optimisation and analyse existing methods with respect to these criteria. Based on this analysis, we propose targeted modifications, such as custom backward passes or adapted maps, that enhance their suitability for analytic gradient-based reinforcement learning. We also extend the applicability of one of the safeguards to state constraints state sets.
@@
We evaluate the approaches in differentiable simulations of various classical control problems. Throughout, we observe sample efficiency and final performance that exceeds or is on par with unsafe training and sampling-based baselines.
@@
  • the first provably safe analytic policy optimisation,
@@
  • and an evaluation on two classical control tasks, demonstrating the potential of provably safe reinforcement learning from analytic gradients.
@@ 2 RELATED WORK
We provide a literature review on the most relevant research areas: analytic gradient-based reinforcement learning, safeguards, and implicit layers.
@@ 2.1 ANALYTIC GRADIENT-BASED REINFORCEMENT LEARNING
Analytic gradient-based reinforcement learning relies on a continuous computational graph from policy actions to rewards, which allows computing the first-order gradient of the reward with respect to the action. Relying on first-order gradient estimators often results in less variance than zeroth-order estimators [GradientOrderDoubt], which are usually obtained using the policy gradient theorem. Less variance leads to faster convergence to local minima of general non-convex smooth objective functions [GradientRL1, GradientRL2]. However, complex environments may lead to optimisation landscapes that are stiff, chaotic or contain discontinuities, which can stifle performance as first-order gradients suffer from empirical bias [GradientOrderDoubt]. Using a smooth surrogate to approximate the underlying noisy reward landscape can alleviate this issue [SHAC]. Moreover, naively backpropagating through time [BPTT] can lead to vanishing or exploding gradients in long trajectories [BPTTProblems].
Various approaches have been introduced to overcome this issue: Policy Optimisation via Differentiable Simulation [PODS] utilises the gradient provided by differentiable simulators in combination with a Hessian approximation to perform policy iteration. It outperforms sampling-based methods. Short-Horizon Actor-Critic (SHAC) [SHAC] tackles the empirical bias of first-order gradient estimators by training a smooth value function through a mean-squared-error loss, with error terms calculated from the sampled short-horizon trajectories through a td-$\lambda$ formulation [RLBible]. It prevents exploding and vanishing gradients by cutting the computational graph deliberately after a fixed number of steps and estimating the terminal value by the critic. The algorithm shows applicability even in contact-rich environments, which tend to lead to stiff dynamics. The successor Adaptive Horizon Actor-Critic [AHAC] has a flexible learning window to avoid stiff dynamics and shows improved performance across the same tasks. Short-Horizon Actor-Critic also inspired Soft Analytic Policy Optimisation [SAPO], which integrates maximum entropy principles to escape local minima.
2.2 SAFEGUARDS
@@
Safeguards are generally categorised according to their safety level [Categorisation, ProvablySafeRL]. Since we seek guarantees, we limit the discussion to hard constraints. Moreover, safeguards must define a differentiable map from unsafe to safe action to allow for backpropagation.
Analytic differentiation and assessing safety necessitate some form of model. However, having the exact model of the system is usually not required. Instead, one can derive conservative approximations of safety based on available data. The two most common techniques to do so are control barrier functions, which guarantee the forward invariance of a set called a safe set [CBF] and reachability analysis, where all possible system behaviours are captured by a reachset-conformant model [Reachability]. Handy safe action sets are robust control invariant sets as they are time-invariant at the cost of potentially conservative results. They can be approximated automatically for nonlinear systems from the state and input constraints of the dynamics [RCI-Sets].
@@
Only maps between continuous action spaces can be differentiable. Krasowski et al. [ProvablySafeRL] present continuous action projection with safe action sets represented by intervals where straightforward re-normalisation is employed to map from the feasible action set. Stolz et al. [RayMasking] generalise this to more expressive sets with their ray mask method. Tabas et al. [Gauge] derive a differentiable bijection based on Minkowski functionals and apply it to power systems. Chen et al. [BoundaryProjectionLoss] define differentiable projection layers relying on convex constraints. While these approaches are, in principle, differentiable, previous work only utilises them to modify policy gradients.
@@ 2.3 IMPLICIT LAYERS
Defining the aforementioned safeguards can often not be done explicitly in closed form. Instead, they can only be formulated implicitly as a separate optimisation problem. Implicit layers [ImplicitLayers, CVXPY-Layers] enable an efficient backpropagation through the solution of this separate optimisation problem without unrolling the solver steps. They decouple the forward and backward pass and analytically differentiate via the implicit function theorem [ImplicitFunctionTheorem], which allows for constant training memory. Implicit layers are a potent paradigm that can be utilised for the tuning of controller parameters [CVXPY-Control], model identification [CVXPY-Models], and safeguarding [BoundaryProjectionLoss]. Given the potential complexity of general optimisation problems being NP-hard, it is crucial to approach the implicit formulation with diligence. If a restriction to convex cone programs is possible, solutions can be computed efficiently in polynomial time [diffcp], thereby facilitating a swift forward pass [GeometricProgramming][ConvexConicPrograms]. Moreover, this allows the utilisation of CVXPY [CVXPY, CVXPY2] to formulate the problem, which automatically picks an efficient solver and translates the problem to the desired solver formulation.
@@ 3 PRELIMINARIES
We use zonotopes to represent safe sets due to their compact representation and closedness under linear maps and Minkowski sums. They are convex, restricted polytopes and are defined as
@@
with centre $c\in\mathbb{R}^{d}$, generator matrix $G\in\mathbb{R}^{d\times n}$, and scaling factors $\beta\in\mathbb{R}^{n}$. Zonotopes with orthogonal generators and $d=n$ are boxes which we use as representations for feasible and noise sets. We utilise the following properties of zonotopes to formulate our safeguards. The Minkowski sum of two zonotopes $\mathcal{Z}_{1},\mathcal{Z}_{2}\subset\mathbb{R}^{d}$ is [RCI-Sets, Eq. 7a]
@@
Translating a zonotope is equivalent to translating the centre. Linearly mapping by $M\in\mathbb{R}^{n\times d}$ yields [RCI-Sets, Eq. 7b]
@@
A support function of a set describes the farthest extent of the set in a given direction. The support function of a zonotope is [SupportFunction, Lemma 1]
@@
A point $p\in\mathbb{R}^{d}$ is contained in a zonotope if [zonotope_point_containment, Eq. 6]
@@
Determining the containment of a zonotope in another zonotope is co-NP complete [zonotope_point_containment] but a sufficient condition for $\mathcal{Z}_{1}\subseteq\mathcal{Z}_{2}$ is [zonotope_containment, Eq. 15]
$1\geq$ $\!\min_{\gamma\in\mathbb{R}^{n_{2}},\Gamma\in\mathbb{R}^{n_{2}\times n_{1}}}$ $\left\lVert\begin{bmatrix}\Gamma&\gamma\end{bmatrix}\right\rVert_{\infty}$ (6a)
@@ 4 PROBLEM STATEMENT
The class of problems we tackle can be described by an augmented Markov decision process $(\mathcal{S},\mathcal{A},\mathcal{A}_{s},P_{f},r)$ of the following elements: the feasible state set $\mathcal{S}\subset\mathbb{R}^{d_{\mathcal{S}}}$, the feasible action set $\mathcal{A}\subset\mathbb{R}^{d}$, the safe action set $\mathcal{A}_{s}\subseteq\mathcal{A}$, the transition probability density function $P_{f}(s_{i+1}|s_{i},a_{i})$, and the reward function $r(s_{i},a_{i},s_{i+1})=r_{i}$. Axis-aligned boxes represent the feasible state and action set and the noise set. For the safeguarding, we employ an overapproximative transition function $f(s_{i},a_{i},w_{i})=s_{i+1}$ with bounded noise $w_{i}\sim\mathcal{W}=\left<c_{\mathcal{W}},G_{\mathcal{W}}\right>\subset\mathbb{R}^{d_{\mathcal{S}}}$, which can, for example, be obtained using reachset-conformant identification [Reachability, Overapproximation, Overapproximation1]. The transition and reward functions are continuously differentiable. The safe action set is either given explicitly as a zonotope, induced by a safe state zonotope $\mathcal{S}_{s}$, or the intersection of both constraints. The induced safe action set is
with the next state set
$\mathcal{S}_{i+1}=f(s_{i},a_{i},\mathcal{W})=\{f(s_{i},a_{i},w_{i})\mid w_{i}\in\mathcal{W}\}\,.$ (8)
We seek a safeguarded, stochastic policy that maximises the expected, discounted, cumulative reward over a finite horizon $N$
$\pi^{*}(a|s)=\!\operatorname*{argmax}_{\pi(a|s)}\operatorname*{\mathbb{E}}_{\begin{subarray}{c}a\sim\pi(a|s)\\ s_{i+1}\sim P_{f}\end{subarray}}\sum_{i=0}^{N}\delta^{i}\,r(s_{i},{a_{s}}_{i},s_{i+1})$ (9)
with the safe action $a_{s}$, the continuously differentiable safeguarding $g_{\mathcal{A}_{s}}(a_{i},s_{i})={a_{s}}_{i}$, and discount factor $\delta\in(0,1)$.

### v2

@@ Abstract
The deployment of autonomous robots in safety-critical applications requires safety guarantees. Provably safe reinforcement learning is an active field of research that aims to provide such guarantees using safeguards. These safeguards should be integrated during training to reduce the sim-to-real gap. While there are several approaches for safeguarding sampling-based reinforcement learning, analytic gradient-based reinforcement learning often achieves superior performance from fewer environment interactions. However, there is no safeguarding approach for this learning paradigm yet. Our work addresses this gap by developing the first effective safeguard for analytic gradient-based reinforcement learning. We analyse existing, differentiable safeguards, adapt them through modified mappings and gradient formulations, and integrate them with a state-of-the-art learning algorithm and a differentiable simulation. Using numerical experiments on three control tasks, we evaluate how different safeguards affect learning. The results demonstrate safeguarded training without compromising performance.
@@ 1 INTRODUCTION
The transfer of physical labour from humans and human-operated machines to robots is a long-standing goal of robotics research. Although robots have been successfully deployed in controlled environments, such as factories, their deployment in human proximity remains challenging [1]. One reason is a lack of safety guarantees to ensure that robots do not harm humans or themselves [2].
@@
A fundamental requirement for safe human-robot interaction is the deployment of controllers with provable safety guarantees. This becomes particularly challenging when using reinforcement learning, which often outperforms classical control methods in uncertain, high-dimensional, and nonlinear environments [3, 4]. To avoid costly and slow real-world training, an agent should preferably train in simulation before deployment to real systems [5, 6]. If the system is safety-critical, applying safeguards already during training is desirable to reduce the sim-to-real gap [7, 8]. Otherwise, the unsafeguarded optimisation may converge to a policy that relies on unsafe states or actions. When the deployment safeguard subsequently restricts the policy, it can become suboptimal or even fail for non-convex objective landscapes, as it has not learned alternative, safe solutions [9]. While crafting reward functions that reliably encode safety requirements could theoretically prevent performance degradation in deployment, this is notoriously difficult without introducing unintended incentives [10, 11]. Moreover, safeguards can aid learning by guiding exploration in challenging solution spaces [12, 13].
In recent years provably safe reinforcement learning has emerged as a research field [14, 15]. Current safeguards are applied in conjunction with reinforcement learning algorithms that rely on the policy gradient theorem to estimate reward landscapes [16, 17, 12, 18, 19, 20]. The advent of differentiable physics simulators [21, 22, 23, 24, 25] allows forgoing this estimation, as a differentiable simulator enables the analytical computation of the reward gradient with respect to actions by differentiating through the dynamics. While these simulators require approximations to remain differentiable, maintaining simulation accuracy is possible. Reinforcement learning algorithms that exploit these gradients promise faster training and better performance [26, 27, 28, 29]. However, existing safeguards for sampling-based reinforcement learning can not be naively applied to these algorithms. Furthermore, there are currently no safeguarding attempts tailored to analytic gradient-based reinforcement learning.
@@
Our work combines state-of-the-art analytic gradient-based reinforcement learning algorithms, differentiable safeguards, and differentiable simulations. As differentiable safeguards, we incorporate a range of provably safe set-based safeguarding methods, whose codomain is a subset of a verified safe action set. This set consists by construction only of safe actions. We formalise desirable safeguarding properties in the context of differentiable optimisation and analyse existing methods with respect to these criteria. Based on this analysis, we propose targeted modifications, such as custom backward passes or adapted maps, that enhance the suitability for analytic gradient-based reinforcement learning. We also extend the applicability of one of the safeguards to state constraints.
@@
We evaluate the provably safe approaches in differentiable simulations of various control problems. We observe sample efficiency and final performance that exceeds or is on par with unsafe training and sampling-based baselines.
@@
  • the first provably safe policy optimisation approach from analytic gradients [^1] Code available at https://github.com/TimWalter/SafeGBPO,
@@
  • and an evaluation on three control tasks, demonstrating the potential of provably safe reinforcement learning from analytic gradients.
@@ 2 RELATED WORK
We provide a literature review on the most relevant research areas: analytic gradient-based reinforcement learning, safeguards, and implicit layers that allow the computation of analytical gradients for optimisation-based safeguards.
@@ 2.1 ANALYTIC GRADIENT-BASED REINFORCEMENT LEARNING
Analytic gradient-based reinforcement learning relies on a continuous computational graph from policy actions to rewards, which allows computing the first-order gradient of the reward with respect to the action via backpropagation. Relying on first-order gradient estimators often results in less variance than zeroth-order estimators [30], which are usually obtained using the policy gradient theorem. Less variance leads to faster convergence to local minima of general non-convex smooth objective functions [26, 27]. However, complex or contact-rich environments may lead to optimisation landscapes that are stiff, chaotic or contain discontinuities, which can stifle performance as first-order gradients suffer from empirical bias [30]. Using a smooth surrogate to approximate the underlying noisy reward landscape can alleviate this issue [29]. Moreover, naively backpropagating through time [31] can lead to vanishing or exploding gradients in long trajectories [32].
[…]
2.2 SAFEGUARDING REINFORCEMENT LEARNING
@@
Safeguards are generally categorised according to their safety level [35, 15]. Since we seek guarantees, we limit the discussion to hard constraints. Moreover, safeguards for analytic gradient-based reinforcement learning must define a differentiable map from unsafe to safe actions to allow for backpropagation.
Within this field of research, two common approaches for enforcing safety guarantees are control barrier functions[36, 37, 38, 39, 40] and set-based reachability analysis [41, 42, 17]. Both necessitate some form of environment model in their basic form, but allow incorporating models identified from data [43, 44, 45]. By finding a control barrier function for a given system, forward invariance of a safe state set can be guaranteed [36]. While mainly used for control-affine systems, solutions for non-affine systems that rely on trainable high-order control barrier functions exist [39]. Nevertheless, finding suitable candidates for control barrier functions for complex systems is nontrivial, and uncertainty handling remains challenging [43]. Therefore, we employ set-based reachability analysis, which employs enclosing models of the true environment dynamics to compute all possible system states [17]. Containment of the reachable state set in a safe state set can be guaranteed by adjusting reinforcement learning actions via constrained optimisation. If robust control invariant sets [46] and reachset-conformant system identification are used [45], this approach can be applied efficiently to non-affine systems with uncertainties.
@@
Differentiable maps between unsafe and safe actions are required to combine the safeguarding approaches with analytic gradient-based reinforcement learning, which are only available for continuous action spaces. Krasowski et al. [15] present continuous action projection with safe action sets represented by intervals, where straightforward re-normalisation is employed to map from the feasible action set. Stolz et al. [12] generalise this to more expressive sets with their ray mask method. Tabas et al. [19] derive a differentiable bijection based on Minkowski functionals and apply it to power systems. Chen et al. [18] define differentiable projection layers relying on convex constraints. Gros et al. [20] define the mapping as an optimisation problem to determine the closest safe action. While these approaches are, in principle, differentiable, previous work only utilises them to modify policy gradients. We utilise and modify boundary projection [20] and ray masking [12] in particular to modify policy behaviour in a differentiable setting.
@@ 2.3 IMPLICIT LAYERS
Defining the safeguards above can often not be done explicitly in closed form. Instead, they can only be formulated implicitly as a separate optimisation problem. Implicit layers [47, 48] enable an efficient backpropagation through the solution of this separate optimisation problem without unrolling the solver steps. They decouple the forward and backward pass and analytically differentiate via the implicit function theorem [49], which allows for constant training memory. Implicit layers are a potent paradigm that can be utilised for the tuning of controller parameters [50], model identification [51], and safeguarding [18]. Given the complexity of general optimisation problems being NP-hard, it is crucial to approach the implicit formulation with diligence. If a restriction to convex cone programs is possible, solutions can be computed efficiently in polynomial time [52], thereby facilitating a swift forward pass [53][54]. Moreover, this allows the utilisation of CVXPY [55, 56] to formulate the problem, which automatically picks an efficient solver and translates the problem to the desired solver formulation.
@@ 3 PRELIMINARIES
We briefly introduce reinforcement learning based on analytical gradients, safe action sets, which serve as the notion of provable safety throughout this work, and the zonotope set representation, which we utilise to describe the safe action sets.
3.1 Analytical gradient-based reinforcement learning
Traditionally, deep reinforcement learning learns an action policy based on scalar rewards without assuming access to a model of the environment dynamics. Prominent algorithms such as REINFORCE [57], Proximal Policy Optimisation [58], or Soft Actor-Critic [59] are based on the policy gradient theorem [60]. This theorem provides an estimator for the gradient of the expected return $J(\theta)$ with respect to the policy parameters $\theta$, given by:
$\frac{\partial}{\partial\theta}J(\theta)=\mathbb{E}_{\pi_{\theta}}\left[\sum_{t=0}^{T}\frac{\partial}{\partial\theta}\log\pi_{\theta}(a_{t}\mid s_{t})Q^{\pi}(s_{t},a_{t})\right]\,,$
where $\pi_{\theta}$ is the parameterised policy, and $Q^{\pi}(s_{t},a_{t})$ denotes the action-value function under policy $\pi$. This gradient estimate can be used to optimise the policy via stochastic gradient ascent.
In contrast, analytical gradient-based reinforcement learning aims to replace this sample-based estimator with a direct gradient computed through a differentiable model of the environment. In such cases, the chain rule can be applied to the full reward computation, yielding an analytical first-order estimate of the policy gradient:
$\frac{\partial}{\partial\theta}J(\theta)=\sum_{t=0}^{T}\left(\frac{\partial r(s_{t},a_{t})}{\partial s_{t}}\frac{\partial s_{t}}{\partial\theta}+\frac{\partial r(s_{t},a_{t})}{\partial a_{t}}\frac{\partial a_{t}}{\partial\theta}\right)$
The term $\frac{\partial s_{t}}{\partial\theta}$ requires backpropagation through time, which can become numerically unstable for long trajectories. This problem motivates the introduction of a regularising critic, resulting in the Short-Horizon Actor-Critic algorithm [29]. For a more detailed review of reinforcement learning methods based on analytical gradients, we refer the interested reader to [29, 34, 25].
3.2 SAFE ACTION SETS
To achieve provable safety, the safety of all traversed states and executed actions must be verifiable. Thus, we introduce a subset of the feasible state set $\mathcal{S}$, the safe state set $\mathcal{S}_{s}\subseteq\mathcal{S}$, containing all states that fulfil all safety specifications. Furthermore, we assume that provable safety is in principle possible, i.e. starting from a safe state, there must exist a sequence of safe actions that ensures the safety of all traversed states
[15, Proposition 1]
$\exists\left(a_{0},a_{1},\dots\right):\mathcal{S}_{i+1}(a_{i},s_{i})\subseteq\mathcal{S}_{s}\quad\forall i\in\mathbb{N}\,\forall s_{0}\in\mathcal{S}_{s}\,,$ (1)
where $\mathcal{S}_{i+1}(a_{i},s_{i})$ denotes the next state set, i.e. the set of reachable states when executing action $a_{i}$ in safe state $s_{i}$. The proposition also implies that at any point in such a sequence, there exists a non-empty safe action set
from which a policy can select actions. In this work, we introduce safeguards $g_{\mathcal{A}_{s}}:\mathcal{X}\mapsto\mathcal{Y}$ with domain $\mathcal{X}\supseteq\mathcal{A}$ and codomain $\mathcal{Y}\subseteq\mathcal{A}_{s}$ that map any feasible, policy-selected action $a_{i}\in\mathcal{A}$ to a safe action $g_{\mathcal{A}_{s}}(a_{i})={a_{s}}_{i}\in\mathcal{A}_{s}$. These safeguards are therefore provably safe by construction.
3.3 ZONOTOPES
We use zonotopes to represent safe sets due to their compact representation and closedness under linear maps and Minkowski sums. Zonotopes are convex, restricted polytopes and are defined as
@@ 3.3 ZONOTOPES
with centre $c\in\mathbb{R}^{d}$, generator matrix $G\in\mathbb{R}^{d\times n}$, and scaling factors $\beta\in{\color[rgb]{0,0,0}[-1,1]}^{n}$. Zonotopes with orthogonal generators and $d=n$ are boxes. We utilise the following properties of zonotopes to formulate our safeguards. The Minkowski sum of two zonotopes $\mathcal{Z}_{1},\mathcal{Z}_{2}\subset\mathbb{R}^{d}$ is [46, Eq. 7a]
@@
Translating a zonotope is equivalent to translating the centre. Linearly mapping by $M\in\mathbb{R}^{{\color[rgb]{0,0,0}m}\times d}$ yields [46, Eq. 7b]
@@
A support function of a set describes the farthest extent of the set in a given direction. The support function of a zonotope in direction $v\in\mathbb{R}^{d}$ is [61, Lemma 1]
@@
A point $p\in\mathbb{R}^{d}$ is contained in a zonotope if [62, Eq. 6]
@@
Determining the containment of a zonotope in another zonotope is co-NP complete [62], but a sufficient condition for $\mathcal{Z}_{1}\subseteq\mathcal{Z}_{2}$ is [63, Eq. 15]
1 1$\geq$1 ≥ $\!\min_{\gamma\in\mathbb{R}^{n_{2}},\Gamma\in\mathbb{R}^{n_{2}\times n_{1}}}$ $\left\lVert\begin{bmatrix}\Gamma&\gamma\end{bmatrix}\right\rVert_{\infty}$ (8a)
@@ 4 PROBLEM STATEMENT
Our work considers constrained Markov decision processes $(\mathcal{S},\mathcal{A},P_{f},r,\mathcal{A}_{s})$ of the following elements:
  • a feasible state set $\mathcal{S}\subset\mathbb{R}^{d_{\mathcal{S}}}$,
  • a feasible action set $\mathcal{A}\subset\mathbb{R}^{d}$,
  • a transition distribution $P_{f}(s_{i+1}|s_{i},a_{i})$,
  • a continuously differentiable reward function $r(s_{i},a_{i},s_{i+1})=r_{i}$,
  • and a safe action set $\mathcal{A}_{s}\subseteq\mathcal{A}$.
We seek a safeguarded, stochastic policy that maximises the expected, discounted return over a finite horizon $N$
$\pi^{*}(a|s)=\!\operatorname*{argmax}_{\pi(a|s)}\operatorname*{\mathbb{E}}_{\begin{subarray}{c}a_{i}\sim\pi(a_{i}|s_{i})\\ s_{i+1}\sim P_{f}\end{subarray}}\sum_{i=0}^{N}\delta^{i}\,r(s_{i},{a_{s}}_{i},s_{i+1})$ (9)
with the safe action ${a_{s}}_{i}{\color[rgb]{0,0,0}=g_{\mathcal{A}_{s}}(a_{i})}$, the continuously differentiable safeguard $g_{\mathcal{A}_{s}}\colon\mathcal{A}\to\mathcal{A}_{s}$, and discount factor $\delta\in(0,1{\color[rgb]{0,0,0}]}$.
4.1 ENSURING COMPUTATIONAL TRACTABILITY
We impose additional requirements on the problem to ease the computational burden of safeguarding. This concerns the representation of all sets as closed, convex sets, such as zonotopes.
Safeguarding is computationally cheap when the problem setting explicitly provides such a representation for the safe action set. However, safeguarding might be computationally intractable if the safe action set needs to be constructed from a safe state set, as specified in Equation˜2, even if the safe state set is available as a closed, convex set. In these cases, we assume that the next state set can be derived using only computations adhering to disciplined convex programming [64]. This allows to replace any constraint on a safe action $a_{s_{i}}\in\mathcal{A}_{s}$ by the state constraint $\mathcal{S}_{i+1}(a_{s_{i}},s_{i})\subseteq\mathcal{S}_{s}$.
In practice, the next state set is often enclosed via reachability analysis, leading to a conservative under-approximation of the safe action set in the current state. However, maintaining Equation˜1 requires tight enclosures, which is an active field of research for complex systems [44, 65, 66, 46, 67, 68].
One method we discuss in particular is obtaining safe state zonotopes via robust control invariant sets [46], which guarantee the existence of an invariance-enforcing controller that can keep all future states within the safe set. This is achieved by enclosing the transition distribution at the current state by a linear transition function with a noise zonotope $\mathcal{W}=\left<c_{\mathcal{W}},G_{\mathcal{W}}\right>\subset\mathbb{R}^{d_{\mathcal{S}}}$. Such an enclosure can, for example, be obtained using reachset-conformant identification [44], which bounds the remainder of the Taylor expansion. Such a transition function allows a linear computation of the next state set
$\mathcal{S}_{i+1}(a_{i},s_{i})=Ma_{i}+\left<c+c_{\mathcal{W}},G_{\mathcal{W}}\right>\,,$ (10)
where $c$ is the offset an $M$ the Jacobian of the linearisation.